Today was programmed to be the last on site, so there was lots to be packed in. The last Megacore was completed over night, and the PELAGRAs floats were recovered after 48 hours deployed.
This morning there was a deep CTD and the plan was for a quick turnround and then two more CTDs and a SAPS dip. But the weather has started to turn nasty and the CTD cable needed some running repairs, so the plan has been revised. The 400m SAPS cast has been made, and the 250m CTD for a post-deployment calibration dip for the sensors that were recovered from PAP1 is reaching completion (at 19:00 GMT)
While all the frantic work was going on around the CTD hanger, the lab work on underway samples, and on CTD water has continued.
I have put together a collage of Kerstin working in the fume cupboard, Sara’s filtration rig, and Jacco and Greg’s FRRF (Fast Repetition Rate Fluorometry) kit (below). Tomorrow everyone will be securing their data and samples and writing their sections of the cruise report.